A box contains 5 brown and 4 white socks. A man takes out two socks. The probability that they are of the same color is:
Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

A box contains 5 brown and 4 white socks. A man takes out two socks. The probability that they are of the same color is:
---------
P(2 brown or 2 white) = P(2brown) + P(2white)
----
= 5c2/9C2 + 4C2/9C2
----
= (10 + 6)/36
---
= 4/9
=============
Cheers,
Stan H.
=============

Answer by Edwin McCravy(20086)

A box contains 5 brown and 4 white socks. A man takes out two socks. The probability that they are of the same color is:
P(1st brown AND 2nd brown) OR P(1st white AND 2nd white) =

and since AND is a clue to multiply and OR is a clue to add,

P(1st brown)×P(2nd brown) + P(1st white)×P(2nd white) =

P(1st brown) = 5/9
P(2nd brown, when 1st was brown) = 4/8 
P(1st white) = 4/9
P(2nd white, when 1st was white) = 3/8

P(1st brown)×P(2nd brown) + P(1st white)×P(2nd white) =

            (5/9)(4/8) + (4/9)(3/8)


                  20/72 + 12/72

                     32/72

                      4/9

Edwin